OpenStudy (anonymous):

antidifferentiate 3x-1/x+1 please help

OpenStudy (dumbcow):

sub u = x+1 du = dx

OpenStudy (dumbcow):

\[\int\limits \frac{3(u-1) -1}{u} du = \int\limits \frac{3u-4}{u} du = \int\limits 3 du - 4\int\limits \frac{1}{u} du\]

hero (hero):

Why not simplify first?

OpenStudy (anonymous):

Thanks that really helps.

hero (hero):

\[\frac{3x - 1}{x+1} = \frac{3x + 3 - 4}{x+1} = \frac{3x + 3}{x + 1} - \frac{4}{x + 1} = 3 - \frac{4}{x + 1}\]

hero (hero):

Now anti-differentiate \[\int\left(3 - \frac{4}{x + 1}\right)dx\]

hero (hero):

You can do that to avoid u-sub

OpenStudy (anonymous):

I actually like this way better. Its easier to work out this way.

hero (hero):

Which way?

hero (hero):

I made it look more complicated than it really is

hero (hero):

I was just showing all the steps